OpenStudy (anonymous):

15a^2b+5a^2-10a factor out

OpenStudy (dumbcow):

what do all 3 terms have in common

OpenStudy (anonymous):

i dont know....

OpenStudy (anonymous):

a Television game show has 5 doors, of which the contest must pick 2. beheind 2 of the doors are expenseive cars, and behind the other 3 doors they are consolation prizes. Find the prob that the contest wins exactly 1 car, no cars, or atleast 1 car

OpenStudy (dumbcow):

15a^2b = 5*3*a*a*b 5a^2 = 5*a*a -10a = -5*2*a All 3 have a 5 and an a

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

take away the 5 and the a from each term and put it on outside 5a(3ab + a -2)

OpenStudy (anonymous):

ok?

OpenStudy (dumbcow):

thats it its now factored

OpenStudy (anonymous):

thats it completely factored??

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ohh ok thanks